# Physics-Guided Learning of Meteorological Dynamics for Weather Downscaling and Forecasting

### v1

@@ A.1. Dataset Description
Huadong dataset: The Huadong dataset encompasses a latitude range from $26.8^{\circ}N$ to $42.9^{\circ}N$ and a longitude range from $112.6^{\circ}E$ to $123.7^{\circ}E$. It comprises a grid of $64\times 44$ cells, with each cell having a grid size of 0.25 degrees in both latitude and longitude. Notably, the Huadong dataset also incorporates Digital Elevation Model (DEM) data to represent terrain information. Since the terrain information usually refers to the boundary layers in the meteorology model instead of an individual weather factor in the PDE, for simplicity, we do not use this information in the paper. The HRES and ERA5 data cover the period from January 3, 2020, to April 1, 2022. The detailed weather factor descriptions of Huadong datasets are in Table 6. The scores of the average of variables reported in Table 2 are computed based on all eight factors. Due to space limits, we only report the specific scores for four factors.
@@
The statistics and weather factor descriptions of Ningbo, Ningxia and WeatherBench datasets are in Table 7, Table 9, and Table 8, respectively. In the forecasting experiments, we divide each dataset into train, validation, and test sets using an 8:1:1 ratio in chronological order. The scores of the average of variables reported in Table 10 are computed based on all eight factors. Due to space limits, we only report the specific scores for four factors.
Table 6. The statistics and weather factor descriptions of the Huadong dataset
Statistics Time range 2020.1.3 00:00-2022.4.1 24:00 Temporal resolution 1-hour Number of time steps 19368 Latitude and longitude range $26.8^{\circ}$-$42.9^{\circ}$, $112.6^{\circ}$-$123.7^{\circ}$ Spatial resolution Origin: $0.25^{\circ}$, $2\times:0.5^{\circ}$, $4\times:1.0^{\circ}$ Number of grids Origin: $64\times 44$, 2×:32×222$\times$:32$\times$ 222 × : 32 × 22, 4×:16×114$\times$:16$\times$ 114 × : 16 × 11, Weather Factors 100m U wind component the horizontal speed of air moving towards the east, at a height of 100 meters above the surface of the Earth 100m V wind component the horizontal speed of air moving towards the north, at a height of 100 meters above the surface of the Earth 10m U wind component the horizontal speed of air moving towards the east, at a height of 10 meters above the surface of the Earth 10m V wind component the horizontal speed of air moving towards the north, at a height of 10 meters above the surface of the Earth 2m temperature the temperature of air at 2m above the surface of land, sea, or inland waters Mean sea level pressure the pressure of the atmosphere at the surface of the Earth, adjusted to the height of mean sea level Surface pressure the pressure of the atmosphere at the surface of land, sea, and inland water Total precipitation the accumulated liquid and frozen water, comprising rain and snow, that falls to the Earth’s surface
Table 7. The statistics and weather factor descriptions of the Ningbo dataset
Statistics Time range 2021.1.1 20:00-2021.4.1 20:00 Temporal resolution 1-hour Number of time steps 2880 Latitude and longitude range $28.85^{\circ}$-$30.56^{\circ}$, $120.91^{\circ}$-$122.29^{\circ}$ Spatial resolution 0.03 degree Number of grids 2726 Weather factors 2m temperature the temperature of air at 2 meters above the surface of land, sea, or inland waters Total precipitation the accumulated liquid and frozen water, comprising rain and snow, that falls to the Earth’s surface 2m relative humidity the measure of the amount of moisture or water vapor present in the air compared to the maximum amount of moisture that the air could hold at a specific temperature, at 2 meters above the surface of land, sea, or inland waters Pressure the pressure of the atmosphere at the surface of land, sea, and inland water 10m relative vorticity the rotation of the air at a height of 10 meters above the Earth’s surface, relative to the Earth’s rotation 10m divergence the measure of the expansion or spreading out of air at a height of 10 meters above the Earth’s surface 10m wind speed the speed at which the wind is blowing at a height of 10 meters above the Earth’s surface 10m wind direction the direction from which the wind is coming at a height of 10 meters above the Earth’s surface 100m wind speed the speed at which the wind is blowing at a height of 100 meters above the Earth’s surface 100m wind direction the direction from which the wind is coming at a height of 10 meters above the Earth’s surface
Table 8. The statistics and weather factor descriptions of the WeatherBench dataset
Statistics Time range 2008.1 -2018.12 Temporal resolution 6-hour Number of time steps 87600 Latitude and longitude range $90^{\circ}$-$90^{\circ}$, $180^{\circ}$-$180^{\circ}$ Spatial resolution 5.625 degree Weather factors 2m_temperature (t2m) the temperature of air at 2 meters above the surface of land, sea, or inland waters Temperature (t) the air temperature at different pressure levels throughout the atmosphere. 10m_u_component_of_wind (u10) the zonal (east-west) component of the wind speed at 10 meters above the ground, where positive values indicate winds blowing from the west to the east. 10m_v_component_of_wind (v10) meridional (north-south) component of the wind speed at 10 meters above the ground, where positive values indicate winds blowing from the south to the north. geopotential (z) the potential energy per unit mass of the Earth’s gravity field at a given point in the atmosphere
Table 9. The statistics and weather factor descriptions of the Ningxia dataset
Statistics Time range 2021.1.1 0:00-2022.1.1 0:00 Temporal resolution 1-hour Number of time steps 8760 Latitude and longitude range $34.5^{\circ}$-$42^{\circ}$, $106^{\circ}$-$116^{\circ}$ Spatial resolution 0.25 degree Number of grids 1271 Weather factors 100m U wind component the horizontal speed of air moving towards the east, at a height of 100 meters above the surface of the Earth 100m V wind component the horizontal speed of air moving towards the north, at a height of 100 meters above the surface of the Earth 10m U wind component the horizontal speed of air moving towards the east, at a height of 10 meters above the surface of the Earth 10m V wind component the horizontal speed of air moving towards the north, at a height of 10 meters above the surface of the Earth 2m temperature the temperature of air at 2m above the surface of land, sea, or inland waters Mean sea level pressure the pressure of the atmosphere at the surface of the Earth, adjusted to the height of mean sea level Surface pressure the pressure of the atmosphere at the surface of land, sea, and inland water Total precipitation the accumulated liquid and frozen water, comprising rain and snow, that falls to the Earth’s surface
@@ A.2.5. Downscaling baselines
  • Bicubic interpolation (Keys, 1981): a two-dimensional interpolation technique that uses the values and gradients of the function at surrounding grid points to obtain a smooth and continuous interpolated result.
@@
  • RCAN (Yu et al., 2021a): a deep model based on ResNet that incorporates a global pooling layer for channel attention.
@@
A.3. More model comparison
Here we plot more results on model comparison before and after physics guidance for forecasting ranges from 1 hour to 7 days on different circumstances in Table 7 (on the 10m wind (U)), Table 8 (on the temperature), and Table 9 (on the surface pressure). PhyDL-NWP shows improvement in all cases.
[…]
Figure 6. Model comparison in Ningxia dataset before and after physics guidance (+) for forecasting ranges from 1 hour to 7 days on the 100m wind (U).
Figure 7. Model comparison in Ningxia dataset before and after physics guidance (+) for forecasting ranges from 1 hour to 7 days on the 10m wind (U).
Figure 8. Model comparison in Ningxia dataset before and after physics guidance (+) for forecasting ranges from 1 hour to 7 days on the temperature.
Figure 9. Model comparison in Ningxia dataset before and after physics guidance (+) for forecasting ranges from 1 hour to 7 days on the surface pressure.

### v2

@@ A.1. Dataset Description
Huadong dataset: The Huadong dataset encompasses a latitude range from $26.8^{\circ}N$ to $42.9^{\circ}N$ and a longitude range from $112.6^{\circ}E$ to $123.7^{\circ}E$. It comprises a grid of $64\times 44$ cells, with each cell having a grid size of 0.25 degrees in both latitude and longitude. Notably, the Huadong dataset also incorporates Digital Elevation Model (DEM) data to represent terrain information. Since the terrain information usually refers to the boundary layers in the meteorology model instead of an individual weather factor in the PDE, for simplicity, we do not use this information in the paper. The HRES and ERA5 data cover the period from January 3, 2020, to April 1, 2022. The scores of the average of variables reported in Table 2 are computed based on all eight factors. Due to space limits, we only report the specific scores for four factors.
@@
In the forecasting experiments, we divide each dataset into train, validation, and test sets using an 8:1:1 ratio in chronological order. The scores of the average of variables reported in Table 5 are computed based on all eight factors. Due to space limits, we only report the specific scores for four factors.
@@ A.2.5. Downscaling baselines
  • Bicubic interpolation: a two-dimensional interpolation technique that uses the values and gradients of the function at surrounding grid points to obtain a smooth and continuous interpolated result.
@@
  • RCAN (Yu et al., 2021): a deep model based on ResNet that incorporates a global pooling layer for channel attention.